OpenStudy (ladylauren):

y = 8x − 9 y = 4x − 1 Part A: Explain how you will solve the pair of equations by substitution or elimination. Show all the steps and write the solution.

OpenStudy (mhchen):

By elimination, that means subtracting the equations together so it would be: y = 8x - 9 - y = 4x - 1 ______________ 0 = 4x - 8 <-- Then you can solve for x By substitution, since y = 4x - 1, you can plug that into the first equation like: 4x - 1 = 8x - 9 <-- Then you can solve for x

OpenStudy (mathmale):

Note that the solution alone is insufficient here; you must explain in words what you must do to solve the system of equations.

OpenStudy (mathmale):

Name the method you intend to use.

OpenStudy (ladylauren):

i will use substitution.

OpenStudy (ladylauren):

8x-9=4x-1 4x-9=-1 x=2